OpenStudy (anonymous):

4. f(x) = tan(x) – 3 The domain is all real numbers and the range is all real numbers? Am i correct or wrong?

OpenStudy (anonymous):

For range, you are correct! Nice job!! For domain, please notice that this is tangent, not sine or cosine. For tangent we have vertical asymptotes in the graph.

OpenStudy (anonymous):

can you draw the graph for me so i can get a better view ?

OpenStudy (anonymous):

okay so what is the domain?

OpenStudy (anonymous):

Use wolframalpha for your graphs I don't think anyone is going to draw the graph for you

OpenStudy (mathmate):

tan(x)-3 has the same domain as tan(x), with which you are probably familiar.

OpenStudy (anonymous):

\[\tan(x)=\frac{\sin(x)}{\cos(x)}\] and so as you may not divide by 0, the domain of tangent is all real numbers except those for which \[\cos(x)=0\]

OpenStudy (anonymous):

of course you want an answer in terms of x and not of cos(x) so you could say all real numbers except \[x=n\pi+\frac{\pi}{2}, n \in \mathbb Z\]